OpenStudy (anonymous):

Algebra 2 help. Medal!

OpenStudy (anonymous):

??

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

heard of `discriminant` before ?

OpenStudy (anonymous):

I don't know exactly what it is but I know the formula -b-4ac

ganeshie8 (ganeshie8):

\(\large b^2 - 4ac\) is called the discriminant

ganeshie8 (ganeshie8):

start by finding the discriminant of first function : \(\large f(x) = x^2+6x-16\)

OpenStudy (anonymous):

ok so 6^2-4(1)(-16)

ganeshie8 (ganeshie8):

\(\large a = 1\) \(\large b = 6\) \(\large c = -16\) discriminant =\(\large b^2-4ac = ?\)

ganeshie8 (ganeshie8):

Yes ! simplify

OpenStudy (anonymous):

is it plus or minus 64

OpenStudy (anonymous):

its either -28 or 100 but I don't know which one

OpenStudy (anonymous):

im pretty sure its 100

ganeshie8 (ganeshie8):

6^2-4(1)(-16) 36 + 64 100

ganeshie8 (ganeshie8):

You're right !

OpenStudy (anonymous):

sweet now how do I figure out if its 2 irrational solutions or 2 rational solutions

ganeshie8 (ganeshie8):

and since the dicriminant is a perfect square(100 = 10^2), f(x) must have `Two real rationals solutions`

OpenStudy (anonymous):

ohhhhh ok thank you! so what about the second equation

ganeshie8 (ganeshie8):

find the discriminant for second equation also

OpenStudy (anonymous):

it would be 6^2-4(1)(1) so the next one would be 32

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

but im still confused with the conditions...

ganeshie8 (ganeshie8):

and since the dicriminant is > 0, but not a perfect square), g(x) must have `Two real irrational solutions`

OpenStudy (anonymous):

ohhhh yes I get it now!!! :) :) :)

ganeshie8 (ganeshie8):

Here are the conditions : discriminant > 0, and a perfect square : two different real RATIONAL solutions discriminant > 0, and NOT a perfect square : two different real IRRATIONAL solutions discriminant = 0 : two same real rational solutions discriminant < 0 : two complex solutions

ganeshie8 (ganeshie8):

good :)

OpenStudy (anonymous):

ok thank you! can you help with a couple more?

ganeshie8 (ganeshie8):

I'll try... ask them in a separate post :)

OpenStudy (anonymous):

ok hang on